Grade 12Mechanics

If the angle between vector sand b is an acute angle then the difference a_b is

Lokmanjari
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
| a-b |= sqrt (a² +b² - 2 ab cos @)
As @ is acute angle
Henceif @>0 but @
| a-b | > |a| - |b|
& if @ = 0°
|a-b| = |a| - |b|